Measure theory

1)find f+ and f- if fx =cosx+1/2 0<x<2π

2)find the measure of the set{x:sinx>=1/2} for 0<x<2π


1
Expert's answer
2021-09-13T14:35:33-0400

1) Consider the inequality "\cos x+1/2>0" for "0<x\leq \pi". On this interval cosine is a monotonely decreasing function, "\cos x>-1/2=\cos(2\pi/3)" iff "0<x<2\pi/3".

Since "\cos(2\pi-x)=\cos x", for "x\in(\pi,2\pi)" "\cos x>-1/2" iff "0<2\pi-x<2\pi/3", or "4\pi/3<x<2\pi".

Finally, for "x\in(0,2\pi)",

if "x\notin(2\pi/3,4\pi/3)" then "(\cos x+1/2)_+ =\cos x+1/2" and "(\cos x+1/2)_- =0".

if "x\in(2\pi/3,4\pi/3)" then "(\cos x+1/2)_- =\cos x+1/2" and "(\cos x+1/2)_+ =0".


2) "\{x\in(0,2\pi):\sin x\geq 1/2\}=(\pi/6, 5\pi/6)", therefore,

"\mu(\{x\in(0,2\pi):\sin x\geq 1/2\})=\mu((\frac{\pi}{6}, \frac{5\pi}{6}))=\frac{5\pi}{6}-\frac{\pi}{6}=\frac{2\pi}{3}"